OpenStudy (anonymous):

translate into an equation. 70 is the sum of 22 and Carlos's savings. use the variable C to represent Carlos's savings.

OpenStudy (anonymous):

@LynFran would it be c+22=70

OpenStudy (steve816):

That is right :)

OpenStudy (anonymous):

i have another problem to solve

OpenStudy (anonymous):

Carmen is saving money to buy a game. so far she has saved $36,which is three-fourths of the total cost of the game. how much does the game cost?

OpenStudy (anonymous):

@steve816

OpenStudy (steve816):

Medal pls?

OpenStudy (shaik0124):

x *(3/4)=36 x*3=36*4 now divide both sides by 3 u will get the game cost